como se resuelve este problema: (x+1)(2x+5)=(2x+3)(x-4)+5

Answer by richwmiller(17219)

2x^2+7x+5 = 2x^2-5x-12+5
2x^2+7x+5 = 2x^2-5 x-7
12x=-12
x=-1
or
(x+1) (2x+5) = (x+1) (2 x-7)
(2x+5) = (2 x-7)
12x=-12
x=-1